OpenStudy (anonymous):

Find the range of the quadratic function. f(x) = (x + 8)2 - 7 A. [-7, ∞) B. (-∞, -8] C. [-8, ∞) D. (-∞, -7]

OpenStudy (anonymous):

its A right?

OpenStudy (jack1):

are u sure that's a quadratic...?

OpenStudy (anonymous):

thats the question i received in my quiz

OpenStudy (jack1):

\[f(x) = (x+8)^{2} -7\] is this ur equation?

OpenStudy (anonymous):

yep

OpenStudy (jack1):

ok then, we can work with that just a heads up though, you can indicate a power by using the ^ symbol, usually located around the 6 on the keyboard

OpenStudy (anonymous):

the anwer is A See , the square on the bracket will always give us a positive ,so the least number you can have in the range is -7 whateve number you add (i.e x=all numbers) range[-7,infn)

OpenStudy (anonymous):

http://prntscr.com/15hgfi

OpenStudy (anonymous):

is it C

OpenStudy (jack1):

a, your y intercept is +1

OpenStudy (anonymous):

so A

OpenStudy (jack1):

yep, A

OpenStudy (anonymous):

You have 332 feet of fencing to enclose a rectangular region. What is the maximum area? A. 6889 square feet B. 6885 square feet C. 110,224 square feet D. 27,556 square feet

OpenStudy (jack1):

yep, sure

OpenStudy (anonymous):

i calculate it and got A as my answer

OpenStudy (jack1):

how did u calculate it?

OpenStudy (anonymous):

P= 2(l+b)

OpenStudy (anonymous):

using the A=lb

OpenStudy (jack1):

2L + 2W = 332 and L x W = y so from eqn 1, L = 332/2 + W so (332/2 + W) x W = y 166W + W^2 = y when derivative of y = 0 = maximum area y' = 2W + 166 0 = 2W +166 -166 = 2W W = 83 therefore as 2L +2W =332 and w = 83, L must =83 also so L x W =A 83 x 83 = 6889 A